OpenStudy (anonymous):

One of the electron transitions in a hydrogen atom produces infrared light with a wavelength of 7.464 x 10^-6 m. What amount of energy causes this transition?

OpenStudy (anonymous):

Try this formula: \(E=\frac{hc}{\lambda}=h\nu\). h is the Planck's constant, \(6.626\times 10^{-34}\). c is speed of light, \(3\times 10^8\). \(\lambda\) is the wavelength. \(\nu\) is the frequency. In this case, try the first one.